What's the value of N? log7n = 2/3log78 I have no idea what I'm even supposed to do first. By the way 2/3 is the fraction two thirds. I'd really appreciate it if you walk me through it. THANKS
Answer by AnlytcPhil(1810)

What's the value of N? log7n = 2/3log78 I have no idea what I'm even supposed
to do first. By the way 2/3 is the fraction two thirds. I'd really appreciate
it if you walk me through it. THANKS
`
log7n = 2/3·log78
`
Use the rule N·logBU = logBUN on the right side
`
log7n = log782/3
`
Now write 8 as 23
`
log7n = log7(23)2/3
`
Now simplify (23)2/3 to 23·2/3 = 21 = 2
`
log7n = log72
`
Now raise both sides to the 7 power, which amounts to dropping the logs
`
n = 2
`
Edwin
AnlytcPhil@aol.com